# Multi-scale Latent Point Consistency Models for 3D Shape Generation

### v2

@@ Abstract
Consistency Models (CMs) have significantly accelerated the sampling process in diffusion models, yielding impressive results in synthesizing high-resolution images. To explore and extend these advancements to point-cloud-based 3D shape generation, we propose a novel Multi-scale Latent Point Consistency Model (MLPCM). Our MLPCM follows a latent diffusion framework and introduces a hierarchy of latent representations, ranging from point-level to super-point levels, each corresponding to a different spatial resolution. We design a multi-scale latent integration module along with 3D spatial attention to effectively denoise the point-level latent representations conditioned on those from multiple super-point levels. Additionally, we propose a latent consistency model, learned through consistency distillation, that compresses the prior into a one-step generator. This significantly improves sampling efficiency while preserving the performance of the original teacher model. Extensive experiments on standard benchmarks ShapeNet and ShapeNet-Vol demonstrate that MLPCM achieves a 100x speedup in the generation process, while surpassing state-of-the-art diffusion models in terms of both shape quality and diversity.
@@ I Introduction
Generative modeling of 3D shapes plays a crucial role in various applications in 3D computer vision and graphics, empowering digital artists to create realistic and high-quality shapes [1, 2, 3, 4, 5, 6, 7]. For these models to be practically effective, they must provide flexibility for interactive refinement, support the synthesis of diverse shape variations, and generate smooth meshes that seamlessly integrate into standard graphics pipelines.
@@
With the rapid advances in generative models for text, images, and videos, significant progress has also been made in 3D shape generation. Techniques based on variational autoencoders (VAEs) [8, 9, 10], generative adversarial networks (GANs)  [11, 12, 13, 14], and normalizing flow models [2, 15, 16] have been proposed to generate 3D shapes. Recently, diffusion-based models [17] and their latent diffusion variants [3, 18] have achieved state-of-the-art results in this domain.
@@
Despite notable progress, substantial obstacles remain when applying diffusion models to 3D shapes that are represented as point clouds. Point sets are unordered, irregular, and often highly non-uniform in density, which complicates both neighborhood construction and feature aggregation [19, 20]. A single radius may fail to capture thin structures while a large radius may blur sharp edges, and k-nearest-neighbor graphs can become unstable under varying sampling density [20, 21]. Real data frequently contain noise, outliers, holes, and partial coverage from limited viewpoints [22, 23], so the denoiser must be robust to missing regions while still reconstructing fine details. In addition, 3D geometry introduces invariances and symmetries that are easy to violate. Models that are not carefully designed with respect to translation, rotation, and scale may overfit to coordinate frames or camera poses [24, 25]. These factors push the denoising network to learn representations that respect both local geometric primitives such as edges, corners, and small parts, and global object level regularities such as symmetry, topology, and part arrangement. Capturing these patterns jointly is essential for producing high quality point sets that are faithful at a small scale while remaining structurally coherent at a large scale.
@@
A promising strategy is to diffuse in a learned latent space where shapes are summarized into compact codes that suppress nuisance variation and concentrate information about structure [26, 3]. Learning in latent space reduces computational cost and can improve stability, yet relying on a single level of latent representation has proved insufficient. A single scale cannot simultaneously preserve tiny geometric details and encode global layout, which leads to oversmoothed reconstructions or broken long range dependencies [27, 28]. Effective 3D generation therefore requires multiple levels or resolutions together with mechanisms that let information flow across them so that coarse latents guide overall structure while finer latents supply detail where needed [3]. A second challenge concerns practicality. The sampling process in diffusion models typically requires many iterative evaluations of the denoiser, which results in slow generation and high memory consumption [29]. Classifier free guidance and other conditioning signals further increase the computational burden [30]. Without acceleration, these costs hinder real world use in interactive modeling, editing, completion from sparse inputs, or deployment within larger 3D pipelines [31, 32]. Reducing the number of steps while preserving quality, for example through consistency distillation or higher order solvers in latent space [33, 34], remains critical for making diffusion-based point cloud generators both accurate and efficient.
@@
  • Extensive experiments show that our method outperforms exiting approaches on the widely used ShapeNet benchmark in terms of shape quality and diversity. Moreover, after consistency distillation, our model is 100x faster than the previous state-of-the-art in sampling.
@@ II-A Diffusion Models
Diffusion models have achieved state-of-the-art results in image synthesis by learning to reverse a corruption process that gradually adds noise to data. In the common discrete-time formulation, the forward process is a fixed Markov chain that diffuses a clean sample into near-Gaussian noise, while the model is trained to predict either the noise or the score of the noised sample so that it can iteratively denoise at test time [29]. A complementary continuous-time view casts diffusion as stochastic differential equations and learns the score function of perturbed data, which unifies score-based generative modeling with diffusion under a single framework and enables alternative samplers [35, 36]. Maximum likelihood training for continuous-time score models further improves statistical efficiency and connects diffusion training to classic likelihood principles [37].
Scaling laws and architectural choices have made diffusion a robust default for high-fidelity image generation. Hierarchical or two-stage pipelines that combine powerful text encoders with diffusion backbones yield strong semantic control and photorealism [38, 39]. Learning the generative process in a compressed latent space rather than pixel space reduces computational burden and memory cost while maintaining visual quality, which has been key to high-resolution synthesis [26]. Compared with variational autoencoders and generative adversarial networks, diffusion models typically provide more stable optimization, reliable likelihood estimation, and improved mode coverage, mitigating common issues such as posterior collapse in VAEs and mode dropping in GANs [40, 41, 42]. During inference, samples are produced by integrating the learned reverse dynamics from noise toward data, which can be interpreted either as reversing a discrete diffusion chain or integrating an ODE or SDE defined by the learned score [29, 36].
@@ II-B Accelerating Diffusion Models
Despite their success, diffusion models are constrained by slow sampling due to many iterative denoising steps. Training-free acceleration exploits the continuous-time formulation to integrate probability-flow ODE trajectories in fewer steps, including high-order solvers such as DPM-Solver and its improved variants, as well as unified predictor–corrector samplers and adaptive step-size schemes [32, 43, 44]. Training-based acceleration modifies objectives or adds light retraining to shorten trajectories: optimized discretization learns sampling times tailored to a fixed compute budget [45], truncated diffusion shortens the denoising horizon [46, 47], and neural-operator style learners amortize parts of the reverse dynamics [48]. Distillation compresses multi-step teachers into fast students by matching trajectories or terminal outputs, yielding few-step or even single-step generators [49, 50].
Recent works push one-step and few-step generation further. Consistency Models enforce a mapping that is consistent along time and can be trained either by distillation or from scratch, enabling rapid one-step sampling [33]. Follow-up advances improve training stability and data-only training, reduce reliance on perceptual losses, and formalize the theory of consistency training [51, 52]. Stochastic consistency distillation introduces noise-aware consistency to better align with the teacher’s trajectories and improve sample quality at low NFE [53]. In parallel, rectified-flow training has been strengthened to compete with distillation in the low-step regime, showing that a single reflow iteration with improved timestep curricula and robust pre-metrics is often sufficient [54]. Flow-matching and flow trajectory distillation style approaches provide another route to few-step or one-step sampling by directly parameterizing ODE paths from noise to data [55, 56]. There is also growing interest in faster training of diffusion and consistency models themselves, which leverages cross-init consistency phenomena and curriculum-style schedules to cut training cost while retaining generation quality [57]. Together, these strategies reduce latency by one to two orders of magnitude, making diffusion practical for interactive applications and deployment under strict compute budgets.
@@
This task is also referred to as shape generation. Achlioptas et al. [58] first proposed to study this problem, developing a straightforward GAN with multiple MLPs in both the generator and discriminator. They also introduced several metrics to evaluate the quality of 3D GANs. Valsesia et al. [59] improved the generator by incorporating graph convolution operations. Liu et al. [60] employed a tree-structured graph convolution network (GCN) to capture the hierarchical information of parent nodes. Gal et al. [61] used a GAN with multiple roots to generate point sets that achieve unsupervised part disentanglement.
Beyond GAN-based methods, there are also approaches rooted in probability theory [62, 63, 64]. ShapeGF [65] used stochastic gradient ascent on an unnormalized probability density to move randomly sampled points toward high-density regions near the surface of a specific shape. DPM [1] conceptualized the transformation of points from a noise distribution to a point cloud as the inverse process of particle diffusion in a thermal system in contact with a heat bath. They modeled this inverse diffusion as a Markov chain conditioned on a particular shape. Yang et al. [2] introduced PointFlow, which generates point sets by modeling them as a distribution of distributions within a probabilistic framework. Kimura et al. [66] presented a flow-based model called ChartPointFlow, which constructs a map conditioned on a label, preserving the shape’s topological structure. Li et al. [67] developed a modified VAE for parts-aware editing and unsupervised point cloud generation.
Several methods have shown their effectiveness in the auto-encoding task using AE architectures, such as l-GAN [58], ShapeGF [65], and DPM [1]. However, despite advancements in point cloud generation, the implicit relationship between GANs and AEs, which offers significant prior information gain, remains largely unexplored.
@@ III-A Hierarchical VAE Framework
where $q_{\phi}(Z^{L}|X)$ and $q_{\phi}(Z^{l}|Z^{l+1},X)$ are assumed to be factorized Gaussian distributions with learnable mean and fixed variance. Our decoder and prior take the simple form $p_{\psi}(X|\mathcal{Z}):=p_{\psi}(X|Z^{0})$ and $p(\mathcal{Z}):=\prod_{l=0}^{L}p(Z^{l})$ respectively. Here $p(Z^{l}):=\mathcal{N}(\bm{0},I)$. The decoder $p_{\psi}(X|Z^{0})$ is parameterized as a factorized Laplace distribution with learnable means and unit scale, corresponding to an L1 reconstruction error.
The overall architecture is illustrated in Figure 1. As shown, both the encoder and the decoder consist of multiple Set Abstraction (SA) and Feature Propagation (FP) modules. Each SA module contains point-voxel convolution (PVC) layers [68], followed by a Grouper block, which includes the sampling and grouping layers introduced by [20]. Each FP module consists of a nearest-neighbor interpolation operation, followed by MLPs and PVC layers.
First Stage Training. We train the encoder and the decoder with a fixed prior by maximizing the modified evidence lower bound (ELBO),
@@ III-B Multi-Scale Latent Point Diffusion As Prior
After learning the encoder and the decoder in the first stage, we fix them and train a denoising diffusion based prior in the latent point space. Note that we have a set of latent representations $\mathcal{Z}=\{Z^{0},\dots,Z^{L}\}$ ranging from point to super-point levels output by the hierarchical encoder. It is natural to select the latent variable with the coarsest scale (lowest spatial resolution), i.e., $Z^{L}$, to build a diffusion prior since it would enable efficient diffusion in a low-dimensional latent space. However, as demonstrated in prior work [3], point-level latent variables remain crucial for producing high-quality 3D shapes. We thus focus on the point-level latent variable $Z^{0}$ and design mechanisms to fuse multi-scale information. Specifically, we introduce a multi-scale latent point diffusion prior as below,
@@
where $\mathcal{Z}^{\setminus{0}}=\{Z^{1},\dots,Z^{L}\}$. Here $p_{\theta}(Z^{l})$ is again a factorized standard Normal distribution, whereas $p_{\theta}(Z^{0}|\mathcal{Z}^{\setminus{0}})$ is a diffusion model. This design is crucial for enabling efficient sampling and consistency distillation, as it requires learning only one conditional prior while keeping others fixed. In contrast, previous work such as LION [3] needs to learn multiple priors corresponding to different levels in the hierarchy, which significantly complicates sampling and distillation. Similar to other diffusion models, our model consists of the forward and the reverse processes.
Forward Process. Following diffusion models [29], given initial latent point feature $Z^{0}\sim q_{\phi}(Z^{0}|X)$ output by the encoder, we gradually add noise as follows,
@@
where the expectation is taken with respect to $t\sim U([T])$, $X\sim p_{\text{data}}(X)$, $\mathcal{Z}\sim q_{\phi}(\mathcal{Z}|X)$, $\epsilon\sim\mathcal{N}(\bm{0},I)$. Here $U([T])$ is the uniform distribution over $\{1,2,\dots,T\}$. Following [3], we utilize the mixed score parameterization, which linearly combines the input noisy sample and the noise predicted by the network, resulting in a residual correction that links the input and output of the noise function $\epsilon_{\theta}$. More details are provided in Supplementary Materials.

### v3

@@ Abstract
[…]
@@ I Introduction
Generative 3D shape modeling is fundamental to many 3D vision and graphics applications, enabling digital artists to produce realistic and high-quality assets [luo2021diffusion, zeng2022lion, du2024generative, liu2024one, chen2024sketch2nerf, wang2024hallo3d]. For these models to be practically effective, they must provide flexibility for interactive refinement, support the synthesis of diverse shape variations, and generate smooth meshes that seamlessly integrate into standard graphics pipelines.
@@
Driven by rapid progress in text, images, and video generation models, 3D shape generation has likewise advanced progress. Techniques based on variational autoencoders (VAEs) [tan2018variational, mittal2022autosdf], generative adversarial networks (GANs)  [wu2016learning, shu20193d, hao2021gancraft], and normalizing flow models [yang2019pointflow, klokov2020discrete] have been proposed to generate 3D shapes. Recently, diffusion-based models [mo2023dit] and their latent diffusion variants [zeng2022lion, ren2024xcube] have achieved state-of-the-art results in this domain.
@@
Despite notable progress, substantial obstacles remain when applying diffusion models to 3D shapes that are represented as point clouds. Point sets are unordered, irregular, and often highly non-uniform in density, which complicates both neighborhood construction and feature aggregation [qi2017pointnet, qi2017pointnet++, 9925147]. A single radius may fail to capture thin structures while a large radius may blur sharp edges, and k-nearest-neighbor graphs can become unstable under varying sampling density [qi2017pointnet++, wang2019dynamic]. Real data frequently contain noise, outliers, holes, and partial coverage from limited viewpoints [yuan2018pcn, tchapmi2019topnet], so the denoiser must be robust to missing regions while still reconstructing fine details. In addition, 3D geometry introduces invariances and symmetries that are easy to violate. Models that are not carefully designed with respect to translation, rotation, and scale may overfit to coordinate frames or camera poses [fuchs2020se, satorras2021n]. These factors push the denoising network to learn representations that respect both local geometric primitives such as edges, corners, and small parts, and global object level regularities such as symmetry, topology, and part arrangement. Capturing these patterns jointly is essential for producing high quality point sets that are faithful at a small scale while remaining structurally coherent at a large scale.
@@
A promising strategy is to diffuse in a learned latent space where shapes are summarized into compact codes that suppress nuisance variation and concentrate information about structure [rombach2022high, zeng2022lion]. Learning in latent space reduces computational cost and can improve stability, yet relying on a single level of latent representation has proved insufficient. A single scale cannot simultaneously preserve tiny geometric details and encode global layout, which leads to oversmoothed reconstructions or broken long range dependencies [vahdat2020nvae]. Effective 3D generation therefore requires multiple levels or resolutions together with mechanisms that let information flow across them so that coarse latents guide overall structure while finer latents supply detail where needed [zeng2022lion]. A second challenge concerns practicality. The sampling process in diffusion models typically requires many iterative evaluations of the denoiser, which results in slow generation and high memory consumption [ho2020denoising]. Classifier free guidance and other conditioning signals further increase the computational burden [ho2022classifier]. Without acceleration, these costs hinder real world use in interactive modeling, editing, completion from sparse inputs, or deployment within larger 3D pipelines [song2020denoising, lu2022dpm]. Reducing the number of steps while preserving quality, for example through consistency distillation or higher order solvers in latent space [song2023consistency, lu2022dpm++], remains critical for making diffusion-based point cloud generators both accurate and efficient.
@@
[…]
@@ II-A Diffusion Models
Diffusion models reach state-of-the-art image synthesis by learning to invert a progressive noising process that corrupts data step by step. Under the standard discrete-time setup, the forward diffusion is a fixed Markov chain that progressively corrupts a clean sample into approximately Gaussian noise. The model is then trained to predict the injected noise or the score of the perturbed sample, enabling iterative denoising during sampling [ho2020denoising]. A complementary continuous-time perspective formulates diffusion as an SDE and learns the score of the perturbed data distribution. This view unifies diffusion and score-based generative modeling in a single framework and supports alternative sampling schemes [song2019generative, song2020score]. Maximum likelihood training for continuous-time score models further improves statistical efficiency and connects diffusion training to classic likelihood principles [song2021maximum].
Scaling laws and architectural choices have made diffusion a robust default for high-fidelity image generation. Hierarchical or two-stage pipelines that combine powerful text encoders with diffusion backbones yield strong semantic control and photorealism [ramesh2022hierarchical, nichol2021glide, shao2024monodiffusion]. Training the generative dynamics in a compact latent space rather than directly in pixel space substantially lowers compute and memory requirements while preserving visual fidelity, and has become a key enabler of high-resolution synthesis [rombach2022high]. Compared with variational autoencoders and generative adversarial networks, diffusion models typically provide more stable optimization, reliable likelihood estimation, and improved mode coverage, mitigating common issues such as posterior collapse in VAEs and mode dropping in GANs [kingma2013auto, sohn2015learning, goodfellow2020generative]. During inference, samples are produced by integrating the learned reverse dynamics from noise toward data, which can be interpreted either as reversing a discrete diffusion chain or integrating an ODE or SDE defined by the learned score [ho2020denoising, song2020score].
@@ II-B Accelerating Diffusion Models
Despite their success, diffusion models are constrained by slow sampling due to many iterative denoising steps. Training-free acceleration exploits the continuous-time formulation to integrate probability-flow ODE trajectories in fewer steps, including high-order solvers such as DPM-Solver and its improved variants, as well as unified predictor–corrector samplers and adaptive step-size schemes [lu2022dpm, lu2025dpm, zhao2023unipc]. Training-based acceleration modifies objectives or adds light retraining to shorten trajectories, where optimized discretization learns sampling times tailored to a fixed compute budget [watson2021learning], truncated diffusion shortens the denoising horizon [lyu2022accelerating, zheng2022truncated], and neural-operator style learners amortize parts of the reverse dynamics [zheng2023fast]. Distillation compresses multi-step teachers into fast students by matching trajectories or terminal outputs, yielding few-step or even single-step generators [salimans2022progressive, meng2023distillation].
Recent works push one-step and few-step generation further. Consistency Models enforce a mapping that is consistent along time and can be trained either by distillation or from scratch, enabling rapid one-step sampling [song2023consistency]. Follow-up advances improve training stability and data-only training, reduce reliance on perceptual losses, and formalize the theory of consistency training [song2023improved, geng2024consistency]. Stochastic consistency distillation introduces noise-aware consistency to better align with the teacher’s trajectories and improve sample quality at low NFE [liu2025scott]. In parallel, rectified-flow training has been strengthened to compete with distillation in the low-step regime, showing that a single reflow iteration with improved timestep curricula and robust pre-metrics is often sufficient [lee2024improving]. Flow-matching and flow trajectory distillation style approaches provide another route to few-step or one-step sampling by directly parameterizing ODE paths from noise to data [liu2022flow, liu2023instaflow]. There is also growing interest in faster training of diffusion and consistency models themselves, which leverages cross-init consistency phenomena and curriculum-style schedules to cut training cost while retaining generation quality [xu2024towards]. Together, these strategies reduce latency by one to two orders of magnitude, making diffusion practical for interactive applications and deployment under strict compute budgets.
@@
This setting is commonly framed as 3D point shape and scene generation [yang2019pointflow, zhao2021point, hu2025large, hu2024rangeldm]. Early work by Achlioptas et al. [achlioptas2018learning] formalized the problem with a simple GAN architecture built from stacked MLPs for both the generator and discriminator, and proposed quantitative criteria for assessing the realism of generated point sets. Subsequent studies strengthened the generative backbone by injecting structural inductive biases. Valsesia et al. [valsesia2018learning] introduced graph-convolutional layers to better exploit neighborhood relations, while Liu et al. [liu2018treegan] adopted a tree-structured GCN to encode hierarchical dependencies among points. More recently, Gal et al. [gal2021mrgan] explored a multi-root GAN formulation that can produce point clouds with emergent, unsupervised part-level factorization.
Not all point-cloud generators are GAN-based. A separate stream of work builds on explicit probabilistic formulations [zhang20223dilg, li2023generalized, cheng2023sdfusion]. ShapeGF [cai2020learning], for instance, defines an (unnormalized) density over 3D space and then uses stochastic gradient ascent to nudge randomly seeded points toward high-density regions, which tend to concentrate around the target surface. Diffusion models offer a different interpretation [10907786, 11207156]. DPM [luo2021diffusion] frames generation as running a particle-diffusion process in reverse. They start from noise and apply a shape-conditioned Markov denoising chain, motivated by analogies to thermal diffusion dynamics. There are also flow- and VAE-style approaches. PointFlow [yang2019pointflow] casts a point set as a distribution of distributions and samples via continuous normalizing flows. ChartPointFlow [kimura2021chartpointflow] follows a similar path but introduces label-conditioned charts to better preserve topology. EditVAE [li2022editvae], in turn, modifies the VAE framework to enable part-aware editing while still supporting unsupervised point-cloud generation.
Several lines of work suggest that plain auto-encoding setups are already quite competitive for learning point-cloud representations and reconstructing shapes sucha as l-GAN [achlioptas2018learning], ShapeGF [cai2020learning], and the diffusion-based DPM [luo2021diffusion]. Yet, even with steady progress in 3D generation, the deeper connection between GAN-based modeling and AE-style learning has received little systematic attention.
@@ III-A Hierarchical VAE Framework
where $q_{\phi}(Z^{L}|X)$ and $q_{\phi}(Z^{l}|Z^{l+1},X)$ are assumed to be factorized Gaussian distributions with learnable mean and fixed variance. Our decoder and prior take the simple form $p_{\psi}(X|\mathcal{Z}):=p_{\psi}(X|Z^{0})$ and $p(\mathcal{Z}):=\prod_{l=0}^{L}p(Z^{l})$ respectively. Here $p(Z^{l}):=\mathcal{N}(\bm{0},I)$. We model the decoder $p_{\psi}(X\mid Z^{0})$ as a fully factorized Laplace likelihood, where the location (mean) parameters are predicted by the network and the scale is fixed to 1. Under this choice, maximizing the reconstruction likelihood is equivalent to minimizing an L1 reconstruction loss.
The overall architecture is illustrated in Figure 1. As illustrated, both the encoder and decoder are built from stacked Set Abstraction (SA) and Feature Propagation (FP) modules. Each SA module contains point-voxel convolution (PVC) layers [liu2019point], followed by a Grouper block, which includes the sampling and grouping layers introduced by [qi2017pointnet++]. Each FP module consists of a nearest-neighbor interpolation operation, followed by MLPs and PVC layers.
First Stage Training. We optimize the encoder and decoder pair under a fixed prior by maximizing a modified evidence lower bound (ELBO),
@@ III-B Multi-Scale Latent Point Diffusion As Prior
After learning the encoder and the decoder in the first stage, we fix them and train a denoising diffusion based prior in the latent point space. Note that we have a set of latent representations $\mathcal{Z}=\{Z^{0},\dots,Z^{L}\}$ ranging from point to super-point levels output by the hierarchical encoder. It is natural to select the latent variable with the coarsest scale (lowest spatial resolution), i.e., $Z^{L}$, to build a diffusion prior since it would enable efficient diffusion in a low-dimensional latent space. However, as demonstrated in prior work [zeng2022lion], point-level latent variables remain crucial for producing high-quality 3D shapes. We thus focus on the point-level latent variable $Z^{0}$ and design mechanisms to fuse multi-scale information. Specifically, we introduce a multi-scale latent point diffusion prior as below,
@@
where $\mathcal{Z}^{\setminus{0}}=\{Z^{1},\dots,Z^{L}\}$. Here $p_{\theta}(Z^{l})$ is again a factorized standard Normal distribution, whereas $p_{\theta}(Z^{0}|\mathcal{Z}^{\setminus{0}})$ is a diffusion model. This design is crucial for enabling efficient sampling and consistency distillation, as it requires learning only one conditional prior while keeping others fixed. In contrast, previous work such as LION [zeng2022lion] needs to learn multiple priors corresponding to different levels in the hierarchy, which significantly complicates sampling and distillation. As in standard diffusion frameworks, our approach comprises a forward noising process and a learned reverse denoising process.
Forward Process. Following diffusion models [ho2020denoising], given initial latent point feature $Z^{0}\sim q_{\phi}(Z^{0}|X)$ output by the encoder, we gradually add noise as follows,
@@
where the expectation is taken with respect to $t\sim U([T])$, $X\sim p_{\text{data}}(X)$, $\mathcal{Z}\sim q_{\phi}(\mathcal{Z}|X)$, $\epsilon\sim\mathcal{N}(\bm{0},I)$. Here $U([T])$ represents the uniform distribution over $\{1,2,\dots,T\}$. Following [zeng2022lion], we utilize the mixed score parameterization, which linearly combines the input noisy sample and the noise predicted by the network, resulting in a residual correction that links the input and output of the noise function $\epsilon_{\theta}$.